OpenStudy (anonymous):

Sooo.... anyone know how to Find the inverse equation of the power function y=2x^4 and is it a function??

OpenStudy (unklerhaukus):

solve for x

OpenStudy (anonymous):

So.... Would you divide by the power of 4?

OpenStudy (unklerhaukus):

divide by 2 first

OpenStudy (unklerhaukus):

then take the forth root of both sides \(\sqrt[4]n\)

OpenStudy (anonymous):

noooooo...... to find an inverse of a function you have to swap x and y. therefore your starting equation will be 2y^4=x then rearrange for y as the subject

OpenStudy (unklerhaukus):

but remember ± because is an even root

OpenStudy (unklerhaukus):

oh am i doing this wrong/

OpenStudy (anonymous):

end result will be \[\sqrt[4]{x/2}\]

OpenStudy (unklerhaukus):

±

OpenStudy (anonymous):

Thanks guys! Sill have a bit more problems. Its hard teaching myself so I'm on here often >.<

OpenStudy (anonymous):

And is that a function?

OpenStudy (anonymous):

its cool umm... no i dont think so... i need to draw the graph to prove it but if it does not obey the vertical line rule then it is not a function

OpenStudy (unklerhaukus):

there are multiple values of the inverse for a single input because of the \(\pm\)